OpenStudy (anonymous):

simplify bt reducing the radical 49√y^21z^14

OpenStudy (anonymous):

are you there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

since there are 21 factors of y and 14 factors of z under the radical... you can take pairs of them out... for example in this more simple version root(y^3) = yroot(y) so you can take 10 y's out and 7 z's... but one y is left inside so its... (y^10z^7)root(y)

OpenStudy (anonymous):

\[49\sqrt{y^21z^14} = 49y ^{10} z^7\sqrt{y}\]

OpenStudy (anonymous):

the problem is \[\sqrt[49]{y ^{21}z ^{14}}\]